OpenStudy (anonymous):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle?

OpenStudy (anonymous):

Select one: a. (x - 1)2 + (y + 2)2 = 3 b. (x + 1)2 + (y - 2)2 = 3 c. (x + 1)2 + (y + 2)2 = 9 d. (x + 1)2 + (y - 2)2 = 9

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

The equation for a circle with center (a,b) and radius r is: (x-a)^2+(y-b)^2=r^2

OpenStudy (kc_kennylau):

Plug in the values

OpenStudy (anonymous):

i think i know it has to be between c and d though

OpenStudy (kc_kennylau):

Seriously, why does everyone here say the answer is between blah blah blah and blah blah blah

OpenStudy (kc_kennylau):

There is always only one answer, and all the other ones are wrong...

OpenStudy (anonymous):

because i said it CAN BE BETWEEN B AND C

OpenStudy (kc_kennylau):

go on go on

OpenStudy (anonymous):

hint: i'll give you a demo a circle has center (-3, 4) and radius 2 equation of circle is (x−h)2+(y−k)2=r2 where (h,k) is the center and r is the radius since in my example h is -3 and k is 4 i will have (x−−3)2+(y−4)2=22 ⟹(x+3)2+(y−4)2=4 do you get the idea?